# Seer: Facilitating Structured Reasoning and Explanation via Reinforcement Learning

### v1

@@ Abstract
Elucidating the reasoning process with structured explanations from question to answer is fundamentally crucial, as it significantly enhances the interpretability and trustworthiness of question-answering (QA) systems. However, structured explanations demand models to perform intricate structured reasoning, which poses great challenges. Most existing methods focus on single-step reasoning through supervised learning, ignoring logical dependencies between steps. Meanwhile, existing reinforcement learning (RL)-based methods overlook the structured relationships, impeding RL’s potential in structured reasoning. In this paper, we propose Seer, a novel method that maximizes a structure-based return to facilitate structured reasoning and explanation. Our proposed structure-based return precisely describes the hierarchical and branching structure inherent in structured reasoning, effectively capturing the intricate relationships between states. We also introduce a fine-grained reward function to meticulously delineate diverse reasoning steps. Extensive experiments show that Seer significantly outperforms state-of-the-art methods, achieving an absolute improvement of 6.9% over RL-based methods on EntailmentBank, a 4.4% average improvement on STREET benchmark, and exhibiting outstanding efficiency and cross-dataset generalization performance.
@@ 1 Introduction
Navigating machines to reason and explain the reasoning process from question to answer is a long-term pursuit in the AI community (McCarthy, 1959; Yu et al., 2023). While existing QA systems adeptly furnish brief supporting evidence (Rajani et al., 2019; DeYoung et al., 2020), they often fail to clarify the reasoning process from prior knowledge to the answer. This means that it is difficult to understand how the answer is reached. However, by elucidating the reasoning process of predicted answers, we can observably improve interpretability, trustworthiness, and debuggability in QA systems (Dalvi et al., 2021; Ribeiro et al., 2023). As illustrated in Figure 1, when answering the question "Which natural material is best for making a table?", the reasoning process with structured explanations, such as entailment trees (Dalvi et al., 2021) or reasoning graphs (Ribeiro et al., 2023), can explain why "sturdy wood" is the predicted answer, which in turn increases trustworthiness.
@@
Method Reasoning Type Runtime Return
@@
However, deriving those complex structured explanations poses a great challenge. Previous methods (Dalvi et al., 2021; Tafjord et al., 2021) consider them as linearized sequences and generate the entire reasoning process at once. However, these methods lack controllability and may hallucinate unreliable reasoning steps. To address these concerns, recent studies (Hong et al., 2022; Neves Ribeiro et al., 2022; Yang et al., 2022) decompose structured explanations and focus on single-step reasoning via supervised learning. Nevertheless, this approach may not always yield optimal results as it fails to consider the interdependencies between different steps. FAME (Hong et al., 2023) attempts to compensate for these shortcomings by leveraging Monte-Carlo planning (Kocsis and Szepesvári, 2006). Yet, as illustrated in Table 1, it significantly increases the running time and inadvertently explores numerous ineffective steps. Furthermore, FAME still concentrates on isolated single-step reasoning, which hardly facilitates multi-step reasoning. Conversely, RLET (Liu et al., 2022) endeavors to enhance the multi-step reasoning through RL. However, RLET still defines the return (a.k.a cumulative rewards) in a chained trajectory, thereby overlooking the inherent tree or graph logical structure in structured reasoning (Dalvi et al., 2021; Ribeiro et al., 2023). Such simplifications may impede the potential of RL for structured reasoning. For instance, RLET may allocate different returns for equivalent trajectories (Yu et al., 2023).
To address the above issues, we propose Seer, a novel method that facilitates Structured rEasoning and Explanation via Reinforcement learning. In structured reasoning, we observe that the logical dependencies between different steps no longer follow a chained trajectory but instead adhere to the inherent tree or graph structure. Therefore, we propose the structure-based return to precisely describe a tree or graph logical structure, effectively capturing the complex interdependencies between different steps. Furthermore, we refine the reward function to meticulously delineate diverse reasoning steps, particularly redundant ones, which do not play a role in the final structured explanations. Through experiments in Sec. 5.4, we find that redundant steps represent the exploration in the environment, and appropriate penalization contributes to improved reasoning performance.
@@
$\bullet$ We also propose the structure-based return to address the intricate interdependencies among different reasoning steps, effectively stimulating the potential of RL in structured reasoning.
$\bullet$ Extensive experiments demonstrate that Seer significantly outperforms state-of-the-art methods. It not only facilitates the effectiveness and efficiency in structured reasoning but also exhibits outstanding cross-dataset generalization performance.
@@ 2.1 Explanation for Question Answering
Extensive research has delved into various forms of explainability in QA systems (Thayaparan et al., 2020; Wiegreffe and Marasovic, 2021; Lamm et al., 2021). Different from the free-form texts susceptible to hallucinations (Rajani et al., 2019; Wei et al., 2022) or the rationales that only provide supporting evidence (DeYoung et al., 2020; Valentino et al., 2021), the structured explanations, such as the entailment trees (Dalvi et al., 2021) and reasoning graphs (Ribeiro et al., 2023), offer a novel way to generate explanations with structured reasoning in a tree or graph structure, delineating what and how knowledge can be integrated to reach the answer. Despite the remarkable interpretability, the intricate structured reasoning also poses significant challenges (Yu et al., 2023; Xu et al., 2023).
@@ 2.2 Natural Language Reasoning
Natural language reasoning, a process that integrates multiple knowledge to derive new conclusions, has attracted significant attention (Saha et al., 2020; Tafjord et al., 2021; Sanyal et al., 2022). Among these, the entailment trees and reasoning graphs, which involve structured reasoning and reasoning path generation tasks, present considerable challenges (Yu et al., 2023). Dalvi et al. (2021) attempts to transform the structured reasoning into a linearized sequence to fit encoder-decoder models, which may generate hallucinations and invalid reasoning. To alleviate this issue, recent studies (Neves Ribeiro et al., 2022; Hong et al., 2022; Neves Ribeiro et al., 2022; Hong et al., 2023) perform premises selection and reasoning in a step-by-step manner. Nevertheless, these methods decompose structured reasoning and solely leverage isolated single-step supervision to train models. This approach neglects the interdependencies between steps, which may not always yield optimal results. Therefore, in light of the advancements of RL in various reasoning tasks (Poesia et al., 2021; Le et al., 2022), RLET (Liu et al., 2022) attempts to incorporate RL into the entailment trees. However, it has to enumerate all potential actions, which is evidently unacceptable for practical scenarios. Furthermore, RLET still defines returns in chained trajectories to facilitate multi-step reasoning, which is obviously not suitable for tree/graph-based structured reasoning. In contrast, our Seer showcases superior adaptability to chained, tree-based, and graph-based structured reasoning via the structure-based return, which significantly enhances both the reasoning performance and efficiency.
@@ 3.1 Task Formulation
As illustrated in Figure 1, the input of the task comprises a set of facts $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and a hypothesis $h$. The output of the task is an entailment tree $T$ (or a reasoning graph) [^1] For the sake of brevity, we formalize the task using the entailment tree (Dalvi et al., 2021) as an example, and the reasoning graph (Ribeiro et al., 2023) is similar.. The entailment tree $T$ consists of tree-structured reasoning, whose leaf nodes are selected from the relevant facts ($x_{*}$) and intermediate nodes represent the derived intermediate conclusions ($i_{*}$). We represent the annotated groundtruth entailment tree as $T_{gold}$, with its leaf nodes signifying $X_{gold}$.
Figure 2: Overall framework of our Seer.
@@ 3.2 Overview
Figure 2 illustrates the overall framework of our proposed Seer, which mainly includes trajectory collection and parameter updates. For trajectory collection, we delineate the state update process into two steps: action generation and conclusion reasoning. The trajectory is produced iteratively until the stopping condition is satisfied (Appendix C.1). For parameter updates, the trajectories are converted to $T_{\text{pred}}$ and aligned with intermediate nodes in $T_{\text{gold}}$ based on maximum Jaccard similarity. Please refer to Appendix C.2 for more alignment details. Then, the policy and critic are updated based on the structure-based return. Algorithm 1 outlines our proposed method for further reference.
@@ 3.3 Fine-grained Component of Seer
We define the state $s_{t}=\{h,P_{t},C_{t}\}$ as a combination of the hypothesis $h$, existing reasoning steps $P_{t}$ and candidate sentences $C_{t}$. $P_{t}$ contains the reasoning steps so far. $C_{t}$ is the set of sentences that can be selected as premises. Each sentence in $C_{t}$ is either unused facts or intermediate conclusions $I_{t}$ generated by previous steps, i.e., $C_{t}=\{X\cup I_{t}\backslash U_{t}\}$, where $U_{t}$ is the set of used sentences. For the initial state, $s_{1}=\{h,P_{1}=\varnothing,C_{1}=X\}$. Evidently, the existing reasoning steps $P_{t}$ and candidate sentences $C_{t}$ entirely depend on the previous state $s_{t-1}$ and its actions $a_{t-1}$. Thus, the state updates process adheres to the Markov decision process (Bellman, 1957).
@@
While the action "Reason: <premises>" facilitates reasoning with combinations of premises, it also significantly increases the complexity of the action space. This is because <premises> can be any combination of elements from the candidate set $C$. Previous studies (Liu et al., 2022; Hong et al., 2022) opted to limit combinations to pairwise premises, resulting in $\binom{n}{2}$ potential actions, where $n$ is the size of the set $C$. However, this approach has apparent limitations. Firstly, as candidate sentences $C$ increases, the number of enumerated actions will experience exponential growth. This makes it computationally expensive and impractical for larger problems. Secondly, by restricting combinations to pairs only, it potentially overlooks the interdependencies among multiple premises, which may limit the effectiveness and richness of the derived conclusions.
To address this issue, we adopt a generative model (Raffel et al., 2020) as the policy $\pi$ to effectively manage the action space $\mathcal{A}$. Despite the vastness of the theoretical action space deriving from "Reason: <premises>", we observe that the number of effective actions in practice is considerably smaller. Therefore, instead of enumerating all possible pairs which can be computationally expensive and may overlook complex interdependencies, our policy $\pi$ samples from the current action space to generate the top-$k$ effective actions.
$a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\sim\pi(a|s_{t}),\quad a\in\mathcal{A}(s_{t})$ (1)
where the input is a linearized state $s_{t}$, which is the concatenation of $h$, $P_{t}$, and $C_{t}$. This not only reduces the computational complexity but also effectively manages the premises with complex combinations, rather than being confined to paired premises. Furthermore, it is important to note that the probabilities of these actions may not sum up to 1, as they are a subset of the entire action space. Therefore, we proceed with re-normalization to form an appropriate probability distribution over the top-$k$ actions, and sample from it to select the current action $a_{t}$.
$\pi^{\prime}(a_{t}^{i}|s_{t})=\frac{\pi(a_{t}^{i}|s_{t})}{\sum_{j=1}^{k}\pi(a_{t}^{j}|s_{t})},\quad i=1,...,k$ (2)
$a_{t}\sim\pi^{\prime}(a|s_{t}),\quad a\in\{a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\}$ (3)
Consequently, this policy allows us to explore the action space more efficiently and effectively.
@@
If the action $a_{t}$ is "Reason: <premises>", we invoke the entailment module to derive the intermediate conclusion to obtain the next state. The entailment module is also a generative model with its input being <premises>. Following (Hong et al., 2022; Liu et al., 2022), we fine-tune the entailment model in a supervised manner and freeze the parameters during the reinforcement learning process, as depicted in Figure 2.
@@
To evaluate the correctness of the entailment tree, Dalvi et al. (2021) proposed an alignment algorithm based on Jaccard similarity to align each intermediate node of the predicted tree $T_{\text{pred}}$ with $T_{\text{gold}}$. However, different from the fully supervised learning methods, we observe that during the RL process, the policy may explore some redundant steps, i.e., these steps do not contribute to the reasoning for the final hypothesis. Existing RL-based work (Liu et al., 2022) simply treats redundant steps with the same penalty as erroneous steps. This simplification may adversely affect the learning process. Firstly, it may discourage necessary exploration in the action space. Secondly, it does not provide meticulous feedback to guide the policy towards optimal policy, since it fails to distinguish between innocuous actions (redundant steps) and incorrect actions (erroneous steps).
To address this issue, we propose a fine-grained reward function that takes into account correct steps, erroneous steps, and redundant steps. For a trajectory $\tau$, we assume that the last intermediate conclusion is our predicted hypothesis, due to the policy deems it should End here. Then, based on the entailment relationships between premises and intermediate conclusions, we backtrack to construct the predicted entailment tree $T_{\text{pred}}$. Note that there might be some steps not participating in the entailment tree, which are regarded as redundant steps. We consider steps that perfectly match via the alignment algorithm (Dalvi et al., 2021) as correct steps and regard others as erroneous steps.
@@
To enhance training stability, we introduce the critic that estimates the state-value function $V(s_{t})$. The input of $V(s_{t})$ is a linearized state, and the output is a scalar representing the return starting from state $s_{t}$. Then, we employ one-step temporal difference strategy (Sutton, 1988) to incrementally adjust these estimated values:
$G_{t}=r_{t}+\gamma V(s_{t+1})$ (5)
where $\gamma$ is the discount factor. Equation 5 brightly describes the chained multi-step reasoning. However, in structured reasoning, reasoning steps typically adhere to inherent tree or graph structures (Dalvi et al., 2021; Ribeiro et al., 2023). The chained structure is merely a special case among them. Thus, equation 5 may not effectively capture the intricate logical dependencies between steps in structured reasoning.
Therefore, we propose the structure-based return to accommodate the logical dependencies inherent in such structured reasoning.
$\hat{G}_{t}=r_{t}+\gamma\frac{1}{|\mathcal{P}(s_{t})|}\sum_{s_{j}\in\mathcal{P}(s_{t})}V(s_{j})$ (6)
where $\mathcal{P}(s_{t})$ represents the parent node of state $s_{t}$ in the entailment tree $T_{\text{pred}}$ or reasoning graph. When $s_{t}\notin T_{\text{pred}}$, $\mathcal{P}(s_{t})=s_{t+1}$. It is worth noting that in the entailment tree, each state typically has only one parent node, i.e., $\hat{G}_{t}=r_{t}+\gamma V(\mathcal{P}(s_{t}))$. Therefore, the entailment tree is also a special case of reasoning graphs (Ribeiro et al., 2023). However, our structure-based return (equation 6) can comprehensively adapt to structured reasoning involving chained, tree-based and graph-based structured scenarios. For example, as shown in Figure 6, for equivalent trajectories $s_{1}\rightarrow s_{2}\rightarrow s_{3}\rightarrow s_{4}$ and $s_{2}\rightarrow s_{1}\rightarrow s_{3}\rightarrow s_{4}$, previous method (Liu et al., 2022) would assign different returns for state $s_{1}$ and $s_{2}$, even though they finally represent the same tree (Figure 6). Conversely, our method, by precisely delineating the intricate interdependencies between reasoning steps, consistently allocates the same return to any equivalent trajectories, thereby enhancing both stability and effectiveness.

### v3

@@ Abstract
Elucidating the reasoning process with structured explanations from question to answer is crucial, as it significantly enhances the interpretability, traceability, and trustworthiness of question-answering (QA) systems. However, structured explanations demand models to perform intricately structured reasoning, which poses great challenges. Most existing methods focus on single-step reasoning through supervised learning, ignoring logical dependencies between steps. Moreover, existing reinforcement learning (RL) based methods overlook the structured relationships, underutilizing the potential of RL in structured reasoning. In this paper, we propose Seer, a novel method that maximizes a structure-based return to facilitate structured reasoning and explanation. Our proposed structure-based return precisely describes the hierarchical and branching structure inherent in structured reasoning, effectively capturing the intricate relationships between different reasoning steps. In addition, we introduce a fine-grained reward function to meticulously delineate diverse reasoning steps. Extensive experiments show that Seer significantly outperforms state-of-the-art methods, achieving an absolute improvement of 6.9% over RL-based methods on EntailmentBank, a 4.4% average improvement on STREET benchmark, and exhibiting outstanding efficiency and cross-dataset generalization performance. Our code is available at https://github.com/Chen-GX/SEER.
@@ 1 Introduction
Navigating machines to understand and articulate the thought process from posing a question to arriving at an answer has been a long-term pursuit in the AI community (McCarthy, 1959; Yu et al., 2023). Current QA explainable systems adeptly furnish brief supporting evidence (Rajani et al., 2019; DeYoung et al., 2020). However, they often fail to clarify the reasoning process from prior knowledge to the derived answer. By elucidating the reasoning process of answers generation from the language models, we can greatly improve interpretability, trustworthiness, and debuggability (Dalvi et al., 2021; Ribeiro et al., 2023). As illustrated in Figure 1, when generating answers for the question "Which natural material is best for making a table?", the reasoning process with structured explanations, such as entailment trees (Dalvi et al., 2021) or reasoning graphs (Ribeiro et al., 2023), explains why "sturdy wood" is the best answer.
@@
Method Training Emphasis Runtime Return
@@
Deriving such complex structured explanations poses a great challenge. Previous methods (Dalvi et al., 2021; Tafjord et al., 2021) consider structured explanations as linearized sequences and generate the entire reasoning process in one go. However, these methods lack controllability and may hallucinate unreliable reasoning steps. To address these concerns, recent studies (Hong et al., 2022; Neves Ribeiro et al., 2022; Yang et al., 2022) decompose structured explanations and focus on single-step reasoning via supervised learning. Nevertheless, this kind of approach may not always yield optimal results as they fail to consider the interdependencies between different steps. FAME (Hong et al., 2023) attempts to compensate for these shortcomings by leveraging Monte-Carlo planning (Kocsis and Szepesvári, 2006), which significantly increases the running time and inadvertently explores numerous ineffective steps (as shown in Table 1). Furthermore, FAME still concentrates on isolated single-step reasoning, which lacks support for structured reasoning. As a general framework for solving sequential decision-making problems, reinforcement learning (RL) is employed in RLET (Liu et al., 2022) to enhance multi-step reasoning. However, RLET defines the return (a.k.a. cumulative reward) using the standard chain structure, thus lacking the ability to represent the tree (Dalvi et al., 2021) or graph (Ribeiro et al., 2023) logical structures inherent in structured reasoning. As a result, the potential of RL for structured reasoning is not fully exploited.
To address the above issues, we propose Seer, a novel method that facilitates Structured rEasoning and Explanation via Reinforcement learning. In structured reasoning, we observe that the logical dependencies between different steps no longer follow a chained trajectory but instead adhere to the inherent tree or graph structure. Therefore, we propose the structure-based return to precisely describe a tree or graph logical structure, effectively capturing the complex interdependencies between different steps. Additionally, we refine the reward function to meticulously delineate diverse reasoning steps, specifically targeting redundant ones that do not contribute to the final structured explanations. Through experiments in Sec. 5.4, we find that redundant steps represent the exploration in the environment, and appropriate penalization contributes to improved reasoning performance.
@@
$\bullet$ We propose the structure-based return to address the intricate interdependencies among different reasoning steps, effectively stimulating the potential of RL in structured reasoning.
$\bullet$ We conduct extensive experiments to demonstrate the superiority of Seer over state-of-the-art methods. Our method facilitates the effectiveness and efficiency of structured reasoning and exhibits outstanding cross-dataset generalization performance.
@@ 2.1 Explanation for Question Answering
Extensive research has delved into various forms of interpretability in QA systems (Thayaparan et al., 2020; Wiegreffe and Marasovic, 2021; Lamm et al., 2021; Chen et al., 2023). Different from the free-form texts susceptible to hallucinations (Rajani et al., 2019; Wei et al., 2022) or the rationales that only provide supporting evidence (DeYoung et al., 2020; Valentino et al., 2021), the structured explanations, such as the entailment trees (Dalvi et al., 2021) and reasoning graphs (Ribeiro et al., 2023), offer a novel way to generate explanations. These structured methods utilize tree or graph formats to clearly outline what information is used and how it is combined to reach the answer. Despite the remarkable interpretability, the intricately structured reasoning also poses significant challenges (Yu et al., 2023; Xu et al., 2023).
@@ 2.2 Natural Language Reasoning
Natural language reasoning, a process that integrates multiple knowledge to derive new conclusions, has attracted significant attention (Saha et al., 2020; Tafjord et al., 2021; Sanyal et al., 2022; Chen et al., 2024). Among these, the entailment trees and reasoning graphs, which involve structured reasoning and reasoning path generation tasks, present considerable challenges (Yu et al., 2023). Dalvi et al. (2021) attempt to transform structured reasoning into a linearized sequence to fit generative models, which may generate hallucinations and invalid reasoning. To alleviate this issue, recent studies (Neves Ribeiro et al., 2022; Hong et al., 2022; Neves Ribeiro et al., 2022; Hong et al., 2023) perform premises selection and reasoning in a step-by-step manner. Nevertheless, these methods decompose structured reasoning and solely leverage isolated single-step supervision to train models. This kind of approach neglects the interdependencies between different steps, which may not always yield optimal results. Therefore, in light of the advancements of RL in various reasoning tasks (Poesia et al., 2021; Le et al., 2022), RLET (Liu et al., 2022) attempts to incorporate RL into the entailment trees. However, it has to enumerate all potential actions, which is unacceptable for practical scenarios. Furthermore, RLET still defines returns in chained trajectories to facilitate multi-step reasoning, which is not suitable for tree/graph-based structured reasoning. In contrast, our Seer showcases superior adaptability to chained, tree-based, and graph-based structured reasoning via the structure-based return, which significantly enhances both the reasoning performance and efficiency.
@@ 3.1 Task Formulation
As illustrated in Figure 1, the input of the task comprises a set of facts $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and a hypothesis $h$. The output of the task is the reasoning steps in a structured form, such as an entailment tree $T$ or a reasoning graph [^1] Although the reasoning graph (Ribeiro et al., 2023) is a more general structure, to be consistent with the majority of previous work, we use the entailment tree (Dalvi et al., 2021) as an example to formalize the task and illustrate our method. Our proposed method is also applicable to the task described in the form of a reasoning graph.. The entailment tree $T$ consists of tree-structured reasoning, whose leaf nodes are selected from the relevant facts ($x_{*}$) and intermediate nodes represent the derived intermediate conclusions ($i_{*}$). We represent the annotated ground-truth entailment tree as $T_{gold}$, with its leaf nodes signifying $X_{gold}$.
Figure 2: Overall framework of Seer. For trajectory rollout, action generation (Policy) and conclusion generation (entitlement) are performed alternately. The orange area details the reasoning process from $s_{t}$ to $s_{t+1}$. For policy optimization, the reward module assigns rewards and updates the policy and critic based on tree or graph structures.
@@ 3.2 Overview
We model the structured reasoning as a reinforcement learning (RL) task, the goal of which is to learn the optimal reasoning policy. Figure 2 illustrates the overall framework of Seer, which mainly includes trajectory rollout and policy optimization. For trajectory rollout, we generate trajectories based on the current policy, and each trajectory is produced iteratively until the stopping criteria are satisfied (Appendix C.1). For policy optimization, we assign rewards to the collected trajectories and update both the policy and critic using the structure-based return. Algorithm 1 (Appendix A) outlines our proposed method for further reference.
@@ 3.3 Fine-grained Component of Seer
At reasoning step $t$, we define the state $s_{t}=\{h,P_{t},C_{t}\}$ as a combination of the hypothesis $h$, existing reasoning steps $P_{t}$ and candidate sentences $C_{t}$. $P_{t}$ contains the reasoning steps so far, and $C_{t}$ is the set of sentences that can be selected as premises. Each sentence in $C_{t}$ is either unused facts or intermediate conclusions $I_{t}$ generated by previous steps, i.e., $C_{t}=\{X\cup I_{t}\backslash U_{t}\}$, where $U_{t}$ is the set of used sentences. For the initial state, $s_{1}=\{h,P_{1}=\varnothing,C_{1}=X\}$.
@@
The action type "Reason: <premises>" induces a large action space, since premises can be any combination of sentences from the candidate set $C$. To enumerate the probabilities of all potential actions and then sample an action to execute, previous studies (Liu et al., 2022; Hong et al., 2022) limit combinations to pairwise premises, such that the action space is reduced to $\binom{n}{2}$, where $n$ is the size of the set $C$. However, such a simplification incurs some potential drawbacks. First, as the number of candidate sentences increases, the number of potential actions grows exponentially. This renders them impractical for complex reasoning tasks with limited computational resources. Second, by restricting combinations to pairs only, the interdependencies among multiple premises are ignored, which may limit the effectiveness and richness of the derived conclusions.
To address this issue, we adopt a generative model to represent the policy $\pi$, which can directly sample from the action space $\mathcal{A}(s_{t})$. Using the generative model essentially expands the action space where the combinations of premises can be arbitrary. This enables the policy to extensively explore better actions during RL training, not limited to paired premises. Further, to speed up RL training, we first generate the top-$k$ actions using policy $\pi$:
$a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\sim\pi(a|s_{t}),\quad a\in\mathcal{A}(s_{t}),$ (1)
where the input is a linearized state $s_{t}$ (i.e., the concatenation of $h$, $P_{t}$, and $C_{t}$). Then, we proceed with re-normalization to form an appropriate probability distribution over the top-$k$ actions, and sample from it to select the action $a_{t}$ to be performed in the current reasoning step, that is,
$\pi^{\prime}(a_{t}^{i}|s_{t})=\frac{\pi(a_{t}^{i}|s_{t})}{\sum_{j=1}^{k}\pi(a_{t}^{j}|s_{t})},\quad i=1,...,k,$ (2)
$a_{t}\sim\pi^{\prime}(a|s_{t}),\quad a\in\{a_{t}^{1},a_{t}^{2},...,a_{t}^{k}\}.$ (3)
@@
If the action $a_{t}$ is "Reason: <premises>", we invoke the entailment module to derive the intermediate conclusion to obtain the next state. The entailment module is also a generative model with its input being <premises>. Following Hong et al. (2022); Liu et al. (2022), we fine-tune the entailment model in a supervised manner and freeze the parameters during the reinforcement learning process, as shown in Figure 2.
@@
To evaluate the correctness of the entailment tree, Dalvi et al. (2021) proposed an alignment algorithm based on Jaccard similarity to align each intermediate node of the predicted tree $T_{\text{pred}}$ with $T_{\text{gold}}$. However, different from the fully supervised learning methods, we observe that during the RL process, the policy explores different actions to identify the optimal reasoning process, inevitably attempting some redundant steps that do not contribute to reaching the final hypothesis. Existing RL-based work (Liu et al., 2022) simply treats redundant steps with the same penalty as erroneous steps. This simplification may negatively affect the learning process which discourages necessary exploration in the action space. Furthermore, it lacks detailed feedback to guide the policy toward optimal policy, as it fails to differentiate between innocuous actions (redundant steps) and incorrect actions (erroneous steps).
To this end, we propose a fine-grained reward function that assigns different reward values for correct steps, erroneous steps, and redundant steps, as shown in Equation 4. For a trajectory $\tau$, we assume that the last intermediate conclusion is our predicted hypothesis since the policy deems it should End here. Then, we backtrack to construct the predicted entailment tree $T_{\text{pred}}$ (see Appendix C.6 for more details). Note that there might be some steps not participating in $T_{\text{pred}}$, which are regarded as redundant steps. Then, as illustrated in Figure 5, we consider steps that perfectly match via the alignment algorithm (Dalvi et al., 2021) as correct steps and regard others as erroneous steps.
@@
To enhance training stability, we introduce the critic to estimate the state-value function $V(s_{t})$. The input of $V(s_{t})$ is a linearized state, and its output is a scalar representing the return (i.e., cumulative reward) when starting from state $s_{t}$. In the simplest case, the return is the chained sum of the rewards. Accordingly, one-step temporal difference (TD) (Sutton, 1988) is often used to estimate $V(s_{t})$, which is updated by the TD-target:
$G_{t}=r_{t}+\gamma V(s_{t+1}),$ (5)
where $\gamma$ is the discount factor. However, in structured reasoning, reasoning steps typically adhere to inherent tree (Dalvi et al., 2021) or graph (Ribeiro et al., 2023) structures, with the chained structure being merely a special case. Thus, Equation 5 just describes the chained multi-step reasoning, which may not effectively capture the intricate logical dependencies between steps in structured reasoning.
Therefore, we propose the structure-based return, where the TD-target is expressed in a more general formulation:
$\hat{G}_{t}=r_{t}+\gamma\frac{1}{|\mathcal{P}(s_{t})|}\sum_{s_{j}\in\mathcal{P}(s_{t})}V(s_{j}),$ (6)
where $\mathcal{P}(s_{t})$ represents the parent node of state $s_{t}$ in the entailment tree $T_{\text{pred}}$ or reasoning graph. When $s_{t}\notin T_{\text{pred}}$, $\mathcal{P}(s_{t})=s_{t+1}$. It can be seen that our structure-based return (Equation 6) adapts to structured reasoning involving chained, tree-based and graph-based structured scenarios. Especially, entailment tree is a special case of the reasoning graph, in which each state typically has only one parent node, and thus Equation 6 degenerates into $\hat{G}_{t}=r_{t}+\gamma V(\mathcal{P}(s_{t}))$. Furthermore, as shown in Figure 6 (Appendix E), for equivalent trajectories $s_{1}\rightarrow s_{2}\rightarrow s_{3}\rightarrow s_{4}$ and $s_{2}\rightarrow s_{1}\rightarrow s_{3}\rightarrow s_{4}$, previous method (Liu et al., 2022) would assign different returns for state $s_{1}$ and $s_{2}$, even though they represent the same tree in the end. Conversely, our method, by precisely delineating the intricate interdependencies between reasoning steps, consistently allocates the same return to any equivalent trajectories, thereby enhancing both stability and effectiveness.